Explain homogeneous and non-homogeneous differential equation and implicit differential equation each with two examples?
1
Expert's answer
2020-10-21T14:57:15-0400

"\displaystyle\textsf{A first order differential equation is said}\\\textsf{to be homogeneous if it may be written as}\\

{\displaystyle f(x,y)\mathrm{d}y=g(x,y)\mathrm{d}x,}\\


\textsf{where}\, f \, \textsf{and}\, g \, \textsf{are homogeneous functions}\\\textsf{of the same degree of}\, x \, \textsf{and}\, y.\\



 \textsf{The functions are homogeneous functions of the}\\\textsf{same degree of}\, x\, \textsf{and}\, y \\ \textsf{if they satisfies the condition} \\

{\displaystyle f(\alpha x,\alpha y)=\alpha ^{k}f(x,y)}\\

{\displaystyle g(\beta x,\beta y)=\beta^{k}g(x,y)}\\
\textsf{For some constant}\, k \, \textsf{and all real numbers}\, \alpha, \beta.\\ \textsf{The constant}\, k\, \textsf{is called the degree of homogeneity.}\\


\textbf{\textsf{Example}}\\

\frac{\mathrm{d}y}{\mathrm{d}x} + 4xy = 0\\

\sin(x)\frac{\mathrm{d}^2y}{\mathrm{d}x^2} + 6\frac{\mathrm{d}y}{\mathrm{d}x} + y = 0\\


\textsf{Nonhomogeneous differential equations are}\\\textsf{the same as homogeneous differential equations,}\\\textsf{except they can have terms involving only},x \\\textsf{(and constants) on the right side,}\\\textsf{as in this equation.}\\

\textsf{The nonhomogeneous differential equations}\\\textsf{is in this format:}\\

y” + p(x)y' + q(x)y = g(x).\\


\textbf{\textsf{Examples}}\\



\frac{\mathrm{d}^2y}{\mathrm{d}x^2} - 4\frac{\mathrm{d}y}{\mathrm{d}x} + 5y = \cos(x)\\


\frac{\mathrm{d}^2y}{\mathrm{d}x^2} - n^2y = 2 + \sin(7x)\\


\textsf{An equation of type}\, F(x, y, y') \, \textsf{where}\, F \, \textsf{is a}\\\textsf{continuous function, is called the}\\\textsf{first order implicit differential equation.}\\


\textsf{If this equation can be solved for}\,y'\\
\textsf{we get one or several explicit}\\\textsf{differential equations of type}\\
y' = f(x, y)\\

\textbf{\textsf{Examples}}\\

9(y')^2 - 4x = 0\\


y = \ln(25 + (y')^2)"